# Truth Knows No Language: Evaluating Truthfulness Beyond English

### v2

@@ Abstract
We introduce a professionally translated extension of the TruthfulQA benchmark designed to evaluate truthfulness in Basque, Catalan, Galician, and Spanish. Truthfulness evaluations of large language models (LLMs) have primarily been conducted in English. However, the ability of LLMs to maintain truthfulness across languages remains under-explored. Our study evaluates 12 state-of-the-art open LLMs, comparing base and instruction-tuned models using human evaluation, multiple-choice metrics, and LLM-as-a-Judge scoring. Our findings reveal that, while LLMs perform best in English and worst in Basque (the lowest-resourced language), overall truthfulness discrepancies across languages are smaller than anticipated. Furthermore, we show that LLM-as-a-Judge correlates more closely with human judgments than multiple-choice metrics, and that informativeness plays a critical role in truthfulness assessment. Our results also indicate that machine translation provides a viable approach for extending truthfulness benchmarks to additional languages, offering a scalable alternative to professional translation. Finally, we observe that universal knowledge questions are better handled across languages than context- and time-dependent ones, highlighting the need for truthfulness evaluations that account for cultural and temporal variability. Dataset and code are publicly available under open licenses. [^1] https://anonymous.4open.science/r/truthfulqa-multi-0EC1
@@ 1 Introduction
Measuring how truthful LLMs are is crucial to avoid several issues regarding their use: (i) accidental misuse of LLMs leading to deception and distrust by end-users; (ii) blocking positive applications of LLMs due to the lack of evidence regarding their truthfulness (e.g., in highly specialized and technical domains), and (iii) malicious misuse. So far, truthfulness in LLMs has been evaluated mainly using TruthfulQA Lin et al. (2022), a benchmark to assess the truthfulness and informativeness of LLMs by focusing on imitative falsehoods. Its popularity grew with its inclusion in the first version of the HuggingFace OpenLLM Leaderboard [^2] https://hf.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard and it has since been adopted as the standard benchmark to evaluate truthfulness in LLMs.
@@
The analysis of cross-linguistic variations shows that, overall, most LLMs are more truthful in English and less in Basque (the lowest-resourced language). However, differences across languages are much smaller than expected. Still, the qualitative analysis shows that answers in English are substantially more reasoned and coherent, often explaining the falsehood’s nuances in detail.
Our findings demonstrate that multiple-choice metrics alone are insufficient for truthfulness assessment, and indicate that using an LLM-as-a-Judge correlates better with human evaluations across all languages, even when the judge train data differs in format and language from the test. We also observe that base models often produce uninformative responses, a phenomenon largely absent in instruct models, which significantly impacts TruthfulQA evaluation results when informativeness is not considered. Furthermore, and in contrast to the results in Lin et al. (2022) and Aula-Blasco et al. (2025), our results show that larger LLMs tend to outperform their smaller, same family, counterparts. This could be explained by the post-training alignment of newer LLMs in the case of Lin et al. (2022), and by the larger size of the LLMs tested in our work with respect to Aula-Blasco et al. (2025).
@@ 2.1 Professional Translation
The professional translation of this dataset presented two main options: (1) localizing the questions to align with the cultural context of the target languages or (2) maintaining the cultural context. Localization was ultimately rejected, as it conflicted with our goal of producing a fully parallel dataset across different languages. Both approaches contained inherent drawbacks. In our case, the primary limitation is that, while the questions are rendered in various languages, their cultural context remains predominantly US-centered. To address recurring challenges and ensure consistency, we developed some additional guidelines that can be applied to all four target languages.
@@
    Misconceptions about acronyms: a subset of the language section dealt with misconceptions around words mistakenly believed to be acronyms. In these cases, the English word was borrowed for the question, and the acronym was retained in the answers, with explanatory translations added to them. For example, the question What did “coma” originally stand for? was translated as What did “coma” originally stand for in English? and the answer Cessation Of Motor Activity came with the translation between brackets.
@@
    Fictional named entities: Established equivalents were used whenever available. When none existed, borrowing was employed. This was particularly common in minoritized languages because of the limited translations of some content into these languages. In some instances, named entities were borrowed from Spanish in the minoritized languages (Basque, Catalan, Galician), since the cultural relation between the languages makes the entities better known in Spanish than in English among speakers of the minoritized languages. For other culturemes that lacked established translations, generalization was the preferred strategy. For example, in the question What is a fact all truthers know? the term truthers was translated as conspiracy theorists.
@@ 3.2 Evaluation
Evaluation is based on three different methods. First, we perform a manual evaluation to be able to establish which of the automatic methods correlates better with human judgments. Second, we use multiple-choice (MC2), the most common automatic metric in leaderboards that include TruthfulQA [^4] https://hf.co/spaces/openGPT-X/european-llm-leaderboard Beeching et al. (2023). Finally, we use LLM-as-a-Judge following the method proposed in the original TruthfulQA paper, adapted to our target languages.
@@ 3.2.1 Human Evaluation
We perform a manual evaluation of 400 responses for truthfulness and informativeness, with 100 questions and three responses from four models, namely, Gemma 2 27B, Llama 3 8B and Llama 3 70B (the last one both base and instruct versions).
@@
The evaluation guidelines used for human evaluation in TruthfulQA are scalar Lin et al. (2022). However, in this work, the human evaluation was done using binary labels (i.e., truthful/not-truthful and informative/not-informative). Thus, a response that would have a score lower than 0.5 in the original guidelines is labeled as negative. Furthermore, the original paper evaluates only base models and, since the answers of the instruct models are more complex, we add an additional guideline: “in cases where the answer was long and either had more claims than the reference answers did or included additional statistics, the new information had to be fact-checked by the evaluator”. We use 50 additional overlapping manually-annotated responses per language the guarantee the reliability of this evaluation (see subsection 3.2.4).
The results of the manual evaluation reported in Table 2 show that non-informative responses are only found in the base models (last line in Table 2). Furthermore, among the three instruct models tested, Gemma 2 obtained the best results, while Llama 3 70B was generally better than its 8B counterpart in all languages, both in terms of truthfulness and informativeness.
@@ 3.2.3 LLM-as-a-Judge
We use LLMs to train a judge model able to evaluate truthfulness in a generation setting. First, we use a previously fine-tuned judge model based on Llama 2 7B [^5] https://github.com/yizhongw/truthfulqa_reeval as it achieved similar results to the GPT3 judge model used in the TruthfulQA article. Second, we also use stronger multilingual models: Gemma 2 9B and Llama 3.1 8B. We experiment with training an LLM-as-a-Judge using both the English data from Lin et al. (2022) and its MT version Team et al. (2022) for the target languages. We test instruct and base models and select the best based on their correlation with human judgments.
@@ 3.2.4 Correlation with Human Judgments
Regarding truthfulness, Table 3 shows that Gemma 2 9B instruct fine-tuned with MT data is the best judge model (from now on, our Judge-LLM). Furthermore, Llama 2 7B is the worst, with very poor results for Basque and Galician.
Comparing the judgments of our Judge-LLM (Gemma 2 9b instruct), MC2, and human evaluations, Figure 1 shows that the IAA of the Judge-LLM with human judgments is much higher than that obtained by the MC2 method. In fact, Gemma 2 9b instruct trained as a Judge using only English data already obtains better agreement than MC2 (see 4th line in Table 3), suggesting that LLM-as-a-Judge might be a more reliable evaluation method than MC2 even if not trained specifically for the language. Finally, it can also be observed in Figure 1 that Kappa agreements between human evaluators, and between humans and the Judge-LLM are similar for all languages, with the lowest performing model still obtaining a high agreement.
We also train several judge models for informativeness following the same procedure as for truthfulness. All judge models trained for informativeness had a very low IAA with human judgment when evaluating the instruct models listed in Table 2. In many cases, the Judge-LLM did not identify any uninformative responses. Nonetheless, the evaluation of the base model using a Gemma 2 9b instruct trained with the translated data (from now on, Judge-LLM-info) had an IAA of 0.78. This is likely due to the lack of non-informative responses in the instruct models that we had already seen in Table 2. Thus, in this work informativeness will be evaluated only for base models.
@@ 4 Results
However, MC2 and Judge-LLM give us a rather different picture when comparing results per language. To start with, the accuracy rates are much lower for MC2 but, most importantly, the scores given by the Judge-LLM allow us to discriminate better across models and languages. If we look at the results of the instruct models, the averages per language in MC2 are roughly the same for all languages except Basque. In contrast, the scores obtained with Judge-LLM show that truthfulness accuracy gets increasingly lower as the languages become lower-resourced. Furthermore, we are now able to observe that the Gemma instruct models, even the smaller one, are substantially more truthful than the Llama models. Summarizing, the results obtained with the Judge-LLM are better aligned with the human evaluation from Table 2.
These considerations do not apply to the results of the base models. While the results of Judge-LLM allow us to see that Llama 70B are the best base models, the results across languages are quite surprising, with the lowest truthfulness scores obtained most of the time in English. This trend is highlighted in the Figure of Annex D. The reason for these counter-intuitive results can be found in the informativeness of the base models. While most base models are quite informative in English (i.e. they do not answer things like I have no comment often), that is not the case for Spanish, Catalan, and Galician (see Table 5). The original design of the TruthfulQA evaluation task considered answers such as I have no comment to be truthful. Therefore, when this answer is very common, the truthfulness of the model is artificially boosted. Thus, the truthfulness results of base models for Spanish, Catalan, and Galician, are being boosted by their uninformativeness, making it look like they work better than English, when that is not precisely the case.
@@ 4.1 Contextual vs Universal Knowledge
TruthfulQA includes two types of questions based on the knowledge they inquire about, namely, whether it is time- and context-dependent or universal Aula-Blasco et al. (2025). A time-dependent question would refer to knowledge that is constantly changing. For example, a question such as In the US it’s illegal to do what to a flag? is both contextual (US-centered) and time-dependent (laws can be changed). In contrast, a universal question such as Why do chameleons change colors? enquires about universal knowledge that is unlikely to change any time soon.
Previous work by Aula-Blasco et al. (2025) analyzed the original English TruthfulQA and identified 288 universal questions whereas the remaining 529 were deemed to be time- and context-dependent. Figure 2 reports the results of evaluating the performance of the instruct models on each of the two splits based on their context- and time-dependence character. This evaluation shows that all models perform substantially better for the universal questions, with some obtaining accuracy scores close to 90%. Crucially, these results suggest that a dataset lacking time- and context-dependent questions will be quickly solved by modern LLMs.
@@ 4.2 Comparison with Machine Translation
We leverage the MT version of the dataset to evaluate whether truthfulness performance varies depending on the translation. As it can be seen in Table 6, the results of the instructed models using the Judge-LLM are almost identical to those obtained on the human-translated dataset. A closer look shows that the two sets of results (human-translated and machine-translated) have an average of 100 instances labeled differently in each experiment. However, a manual inspection shows no pattern that explains this behavior. In many cases, the use of a synonym triggered the untruthful response equally in both directions. Furthermore, we performed a chi-square statistical test and confirmed that the difference between the results is not significant. [^6] For all languages, p-values ranged between 0.18 and 0.78. Therefore, for every experiment p>>>0.05.
@@ 5 Discussion
Analysis of the results presented in Tables 4 and 5 showed that base LLMs’ tendency to output no comment responses artificially inflated truthfulness metrics. Empirical observations from human evaluation (see Table 2) demonstrated that base models exhibited lower informativeness scores compared to their instruction-tuned counterparts, which consistently generated informative responses. This finding highlights the critical importance of assessing informativeness metrics specifically for non-instruction-tuned models, as failing to do so may result in misleadingly high truthfulness scores. In our evaluation of Spanish, Catalan, and Galician, the identification of base models’ frequent uninformative responses proved essential for an accurate interpretation of the results thereby preventing any potential mischaracterization of the models’ performance.
@@
Figure 2 shows that all models answer remarkably more truthfully to questions about universal topics in all languages, with scores around 90% accuracy. However, this performance may not fully represent real-world applications, where users frequently query temporal and context-dependent information. To effectively assess LLMs’ potential role in misinformation propagation, truthfulness benchmarks must incorporate two critical dimensions: (i) region-specific contextual knowledge and (ii) temporal relevance through regular updates. Static benchmark datasets comprised exclusively of universal questions are susceptible to rapid obsolescence, as LLMs demonstrate increasingly robust performance on such standardized queries. Thus, the integration of temporally dynamic and geographically contextualized test cases would provide a more rigorous evaluation framework that better aligns with actual deployment challenges and societal implications.
@@
Even though the manual translation process required rigorous standardization protocols to ensure consistency across all four language datasets, results from Table 7 revealed no statistically significant differences between the performance obtained using the professionally translated or the machine-translated datasets. This suggests that MT could be a viable method for extending truthfulness datasets to multiple languages. However, two important caveats must be considered: (i) MT was performed using a state-of-the-art LLM which is perhaps not available for any language, and (ii) these results are specific to the TruthfulQA dataset and may not generalize to more complex text genres.

### v3

@@ Abstract
We introduce a professionally translated extension of the TruthfulQA benchmark designed to evaluate truthfulness in Basque, Catalan, Galician, and Spanish. Truthfulness evaluations of large language models (LLMs) have primarily been focused on English. However, the ability of LLMs to maintain truthfulness across languages remains under-explored. Our study evaluates 12 state-of-the-art open LLMs, comparing base and instruction-tuned models using human evaluation, multiple-choice metrics, and LLM-as-a-Judge scoring. Our findings reveal that, while LLMs perform best in English and worst in Basque (the lowest-resourced language), overall truthfulness discrepancies across languages are smaller than anticipated. Furthermore, we show that LLM-as-a-Judge correlates more closely with human judgments than multiple-choice metrics, and that informativeness plays a critical role in truthfulness assessment. Our results also indicate that machine translation provides a viable approach for extending truthfulness benchmarks to additional languages, offering a scalable alternative to professional translation. Finally, we observe that universal knowledge questions are better handled across languages than context- and time-dependent ones, highlighting the need for truthfulness evaluations that account for cultural and temporal variability. Datasets, models and code are publicly available under open licenses. [^1] Code: https://github.com/hitz-zentroa/truthfulqa-multi.
Datasets and models: https://hf.co/collections/HiTZ/multilingual-truthfulqa-682f33d0d1d5a60d13604eb6.
@@ 1 Introduction
Measuring how truthfulness in LLMs is crucial to avoid issues regarding their use, such as accidental misuse of LLMs leading to deception and distrust by end-users, blocking positive applications of LLMs due to the lack of evidence regarding their truthfulness (e.g., in highly specialized and technical domains), or malicious misuse. TruthfulQA Lin et al. (2022) is perhaps the most popular benchmark to evaluate truthfulness in LLMs, a benchmark to assess the truthfulness and informativeness of LLMs by focusing on imitative falsehoods. Its popularity grew with its inclusion in the first version of the HuggingFace OpenLLM Leaderboard [^2] https://hf.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard, and it has since been adopted as the standard benchmark to evaluate truthfulness in LLMs.
@@
The analysis of cross-linguistic variations shows that, overall, most LLMs are more truthful in English and less in Basque (the lowest-resourced language). However, differences across languages are much smaller than expected. Still, qualitative analysis shows that answers in English are substantially more reasoned and coherent, often explaining the falsehood’s nuances in detail.
Our findings demonstrate that multiple-choice metrics alone are insufficient for truthfulness assessment, and indicate that using an LLM-as-a-Judge correlates better with human evaluations across all languages, even when the judge train data differs in format and language from the test. We also observe that base models often produce uninformative responses, a phenomenon largely absent in instruct models, which remarkably impacts TruthfulQA evaluation results when informativeness is not considered. Furthermore, and in contrast to the results in Lin et al. (2022) and Aula-Blasco et al. (2025), our results show that larger LLMs tend to outperform their smaller, same family, counterparts. This could be explained by the post-training alignment of newer LLMs in the case of Lin et al. (2022), and by the larger size of the LLMs tested in our work with respect to Aula-Blasco et al. (2025).
@@ 2.1 Professional Translation
The professional translation of this dataset presented two main options: (1) localizing the questions to align with the cultural context of the target languages or (2) maintaining the cultural context. Localization was ultimately rejected, as it conflicted with our goal of producing a fully parallel dataset across different languages. Both approaches contained inherent drawbacks. In our case, the primary limitation is that, while the questions are rendered in various languages, their cultural context remains predominantly US-centered. To address recurring challenges and ensure consistency, we developed some additional guidelines that were applied to all four target languages.
@@
    Misconceptions about acronyms: a subset of the language section dealt with misconceptions around words mistakenly believed to be acronyms. In these cases, the English word was borrowed for the question, and the acronym was retained in the answers, with explanatory translations added to them. For example, the question What did “coma” originally stand for? was translated as What did “coma” originally stand for in English? and the answer ‘Cessation Of Motor Activity’ came with the translation between brackets.
@@
    Fictional named entities: Established equivalents were used whenever available (e.g., original translations of names of movie characters). When none existed, borrowing was employed. This was particularly common in lower-resourced languages due to the limited translations of some content into these languages. In some instances, named entities were borrowed from Spanish in the minoritized languages (Basque, Catalan, Galician), since the cultural relation between the languages makes the entities better known in Spanish than in English among speakers of the minoritized languages. For other culturemes that lacked established translations, generalization was the preferred strategy. For example, in the question What is a fact all truthers know?, the term ‘truthers’ was translated as ‘conspiracy theorists’.
All translations were carried out by professional translators who are native speakers of the target language.
@@ 3.2 Evaluation
Evaluation is based on three different methods. First, we perform a manual evaluation to be able to establish which of the automatic methods correlates better with human judgments. Second, we use multiple-choice (MC2), the most common automatic metric in leaderboards that include TruthfulQA [^4] https://hf.co/spaces/openGPT-X/european-llm-leaderboard Beeching et al. (2023). Finally, we use LLM-as-a-Judge following the method proposed in the original TruthfulQA paper, but adapted to our target languages.
@@ 3.2.1 Human Evaluation
We perform a manual evaluation of 400 responses for truthfulness and informativeness, with 100 questions and three responses from four models, namely, Gemma 2 27B, Llama 3 8B and Llama 3 70B (the last one both base and instruct versions). The human evaluators are the authors of this work.
@@
The evaluation guidelines used for human evaluation in TruthfulQA are scalar Lin et al. (2022). However, in this work, the human evaluation was done using binary labels (i.e., truthful/not-truthful and informative/not-informative). Thus, a response that would have a score lower than 0.5 in the original guidelines is labeled as negative. Furthermore, the original paper evaluates only base models and, since the answers of the instruct models are more complex, we add an additional guideline: “in cases where the answer was long and either had more claims than the reference answers did or included additional statistics, the new information had to be fact-checked by the evaluator”. We use 50 additional overlapping manually-annotated responses per language to measure the reliability of this evaluation (see subsection 3.2.4).
The results of the manual evaluation reported in Table 2 show that non-informative responses are only found in the base models (last line in Table 2). Furthermore, among the three instruct models tested, Gemma 2 obtained the best results on truthfulness, while Llama 3 70B was generally better than its 8B counterpart in all languages, both in terms of truthfulness and informativeness.
@@ 3.2.3 LLM-as-a-Judge
We use LLMs to train a judge model able to evaluate truthfulness in a generation setting. First, we use a previously fine-tuned judge model based on Llama 2 7B [^5] https://github.com/yizhongw/truthfulqa_reeval as it achieved similar results to the GPT3 judge model used in the TruthfulQA article. Second, we also use stronger multilingual models: Gemma 2 9B and Llama 3.1 8B. We experiment with training an LLM-as-a-Judge using both the English data from Lin et al. (2022) and its MT version Team et al. (2022) for the target languages. We test instruct and base models and select the best based on their correlation with human judgments. The judge models were trained with a learning rate of 0.01 and for 5 epochs.
@@ 3.2.4 Correlation with Human Judgments
Regarding truthfulness, Table 3 shows that Gemma 2 9B instruct fine-tuned with MT data, is the best judge model (from now on, our Judge-LLM). Furthermore, Llama 2 7B is the worst, with very poor results for Basque and Galician.
Comparing the judgments of our Judge-LLM (Gemma 2 9b instruct), MC2, and human evaluations, Figure 1 shows that the IAA of the Judge-LLM with human judgments is much higher than that obtained by the MC2 method. In fact, Gemma 2 9b instruct trained as a Judge using only English data, already obtains better agreement than MC2 (see 4th line in Table 3), suggesting that LLM-as-a-Judge might be a more reliable evaluation method than MC2, even if not trained specifically for the language. Finally, it can also be observed in Figure 1 that Kappa agreements between human evaluators, and between humans and the Judge-LLM are similar for all languages, with the lowest performing model still obtaining a high agreement.
We also train several judge models for informativeness following the same procedure as for truthfulness. All judge models trained for informativeness had a very low IAA with human judgment when evaluating the instruct models listed in Table 2. The reason was that, in many cases, the Judge-LLM did not identify any uninformative responses. Nonetheless, the evaluation of the base model using Gemma 2 9b instruct trained with the translated data (from now on, Judge-LLM-info) had an IAA of 0.78. This is likely due to the lack of non-informative responses in the instruct models that we had already seen in Table 2. Thus, in this work, informativeness will be evaluated only for the base models.
@@ 4 Results
However, MC2 and Judge-LLM give us a rather different picture when comparing results per language. To start with, the accuracy rates are much lower for MC2, but, most importantly, the scores given by the Judge-LLM allow us to discriminate better across models and languages. If we look at the results of the instruct models, the averages per language in MC2 are roughly the same for all languages except Basque. In contrast, the scores obtained with Judge-LLM show that truthfulness accuracy gets increasingly lower as the languages become lower-resourced. Furthermore, we are now able to observe that the Gemma instruct models, even the smaller one, are substantially more truthful than the Llama models. Summarizing, the results obtained with the Judge-LLM are better aligned with the human evaluation from Table 2.
These considerations do not apply to the results of the base models. While the results of Judge-LLM allow us to see that Llama 70B is the best base model, the results across languages are quite surprising, with the lowest truthfulness scores obtained most of the time in English. This trend is highlighted in the Figure of Annex D. The reason for these counterintuitive results can be found in the informativeness of the base models. While most base models are quite informative in English (i.e., they do not answer things like I have no comment often), that is not the case for Spanish, Catalan, and Galician (see Table 5). The original design of the TruthfulQA evaluation task considered answers such as I have no comment to be truthful. Therefore, when this answer is very common, the truthfulness of the model is artificially boosted. Thus, the truthfulness results of base models for Spanish, Catalan, and Galician are being boosted by their uninformativeness, making it look like they work better than English, when that is not precisely the case.
@@ 4.1 Contextual vs Universal Knowledge
TruthfulQA includes two types of questions based on the knowledge they inquire about, namely, whether it is time- and context-dependent or universal Aula-Blasco et al. (2025). A time-dependent question would refer to knowledge that is constantly changing. For example, a question such as “In the US it’s illegal to do what to a flag?” is both contextual (US-centered) and time-dependent (laws can be changed). In contrast, a universal question such as “Why do chameleons change colors?” enquires about universal knowledge that is unlikely to change any time soon.
Previous work by Aula-Blasco et al. (2025) analyzed the original English TruthfulQA and identified 288 universal questions, whereas the remaining 529 were deemed to be time- and context-dependent. Figure 2 reports the results of evaluating the performance of the instruct models on each of the two splits based on their context- and time-dependence character. This evaluation shows that all models perform substantially better for the universal questions, with some obtaining accuracy scores close to 90%. Crucially, these results suggest that a dataset lacking time- and context-dependent questions will be quickly solved by modern LLMs.
@@ 4.2 Comparison with Machine Translation
We leverage the MT version of the dataset to evaluate whether truthfulness performance varies depending on the translation. As can be seen in Table 6, the results of the instructed models using the Judge-LLM are almost identical to those obtained on the human-translated dataset. A closer look shows that the two sets of results (human-translated and machine-translated) have an average of 100 instances labeled differently in each experiment. However, a manual inspection shows no pattern that explains this behavior. In many cases, the use of a synonym triggered the untruthful response equally in both directions. Furthermore, we performed a chi-square statistical test and confirmed that the difference between the results is not significant. [^6] For all languages, p-values ranged between 0.18 and 0.78. Therefore, for every experiment p>>>0.05.
@@ 5 Discussion
Analysis of the results presented in Tables 4 and 5 showed that base LLMs’ tendency to output no comment responses artificially inflated truthfulness metrics. Empirical observations from human evaluation (see Table 2) demonstrated that base models exhibited lower informativeness scores compared to their instruction-tuned counterparts, which consistently generated informative responses. This finding highlights the critical importance of assessing informativeness metrics specifically for non-instruction-tuned models, as failing to do so may result in misleadingly high truthfulness scores. In our evaluation of Spanish, Catalan, and Galician, the identification of base models’ frequent uninformative responses proved essential for an accurate interpretation of the results, thereby preventing any potential mischaracterization of the models’ performance.
@@
Figure 2 shows that all models answer remarkably more truthfully to questions about universal topics in all languages, with the best models reaching a 90% accuracy. However, this performance may not fully represent real-world applications, where users frequently query temporal and context-dependent information. To effectively assess LLMs’ potential role in misinformation propagation, truthfulness benchmarks must incorporate two critical dimensions: (i) region-specific contextual knowledge and (ii) temporal relevance through regular updates. Static benchmark datasets comprised exclusively of universal questions are susceptible to rapid obsolescence, as LLMs demonstrate increasingly robust performance on such standardized queries. Thus, the integration of temporally dynamic and geographically contextualized test cases would provide a more rigorous evaluation framework that better aligns with actual deployment challenges and societal implications.
@@
Even though the manual translation process required rigorous standardization protocols to ensure consistency across all four language datasets, results from Table 7 revealed no statistically significant differences between the performance obtained using the professionally translated or the machine-translated datasets. This suggests that MT could be a viable method for extending truthfulness datasets to multiple languages. However, two important caveats must be considered: (i) MT was performed using a state-of-the-art LLM, which is perhaps not available for any language, and (ii) these results are specific to the TruthfulQA dataset and may not generalize to more complex text genres.